henytakeAssighmentMain.dofinvoker assignmentstak or assignment take8inprogress falseAI Product-Costing Accuracy, Plantwide and Departmental Rates, ABC Escuha Company produces two type of calculators: sdentific and business. Both products pass through two producdng departments. The business Calcoulator is by far the most popular. The following data have been gathered for these two products: Product-Related Data Scientific Business Units produced per year Prime costs Direct labor hours Machine hours Production runs Inspection hours Maintenance hours 76,000 $239,000 104,800 52,200 100 2,000 2,250 760,000 $2,390,000 1,048,000 522,000 150 3,000 9,000 Department Data Department 1 Department 2 Direct labor hours: 78,000 125.211 203,760 26,800 922 240 Scientific calculator Business calculator Total Machine hours: 26,500 25,700 421 480 447 190 Scientific calculator 100.520 Business calculator Kpr Previous Check My Work 3 more Check My Work uses remaining for Grading Email Instructor Save and Exit Submit 211 AM ^ョads) ENG 29/01/2019
Total 447,180 127,020 Overhead costs: Setup costs Inspection costs Power Maintenance $225,500 175,000 270,000 212.000 $882,500 $225,500 175,000 160,000 215.500 $276,000 Total Required 1. Compute the overhead cost per unit for each product using a plantwide, unit-based rate using direct labor hours. Round your calculations and answers to the nearest cent. Plantwide rate Overhead cost per unit: Scientific Business 2. Compute the overhead cost per unit for each product using departmental rates. In calculating departmental rates, use machine hours for Department 1 and direct labor hours for Department 2. Round departmental rates and the overhead cost per unit Departmental rates: Department 1 Department 2 per direct labor hour per unit per unit to the nearest cent if rounding is required. per machine hour per direct labor hour Previous Check My Work 3 more Check My Work uses remaining Email Instructor Save and Exit Submit Assignment for Grading 211 AM 29/01/2019
and direct labor hours for Department 2. Round departmental rates and the overhead cost per unit to the nearest cent if rounding is required. Departmental rates: Department 1 Department 2 Overhead cost per unit: Scientific Business Repeat using direct labor hours for Department 1 and machine hours for Department 2. Round departmental rates and the overhead cout pes unit to nearest cent if rounding is required. Departmental rates: Department 1 Department 2 Overhead cost per unit: Scientific Business per machine hour per direct labor hour per unit per unit the per direct labor hour per machine hour per unit per unit 3. Compute the overhead cost per unit for each product using activity-based costing. Round your calculations and answers to the nearest cent. Overhead Cost per unit Der unit Scientific Previous Check My Work 3 more Check My Work uses remaining. Email Instructor Save and Exit Submit Assignment for Gras 212 AM 29/01/20
Answer #1
1) Plant wide rate = (882500+776000)/(203760+949040) = $                1.44 per direct labor hour
Overhead cost per unit: Total overhead
Scientific = (104800*1.44)/76000 = $                1.99 per unit
Business = (1048000*1.44)/760000 = $                1.99 per unit $       16,60,032
2) Departmental rates:
Department 1 = (882500/447180) = $                1.97 per MH
Department 2 = (776000/949040) = $                0.82 per DLH
Overhead cost per unit:
Scientific = (25700*1.97+26800*0.82)/76000 = $                0.96 per unit $       16,59,157
Business = (421480*1.97+922240*0.82)/760000 = $                2.09 per unit
Department 1 = (882500/203760) = $                4.33 per DLH
Deparment 2 = (776000/127020) = $                6.11 per MH
Overhead cost per unit:
Scientific = (78000*4.33+26500*6.11)/76000 = $                6.57 per unit $       16,58,251
Business = (125760*4.33+100500*6.11)/760000 = $                1.52 per unit
3) Activity rates:
Cost pool Overhead Cost Activity Volume Activity Rate
Set up costs 451000 250 $         1,804.00 per set up
Inspection costs 350000 5000 $               70.00 per inspection hour
Power 430000 574200 $                 0.75 per machine hour
Maintenance 427500 11250 $               38.00 per maintenance hour
Total 1658500
Allocation of OH to products:
Cost pool Activity rate Overhead allocated
Scientific Business Total
Set up costs $        1,804.00 $      1,80,400 $         2,70,600 $        4,51,000
Inspection costs $              70.00 $      1,40,000 $         2,10,000 $        3,50,000
Power $                0.75 $          39,150 $         3,91,500 $        4,30,650
Maintenance $              38.00 $          85,500 $         3,42,000 $        4,27,500
Total $      4,45,050 $       12,14,100 $     16,59,150
Overhead cost per unit using ABC:
Scientific = 445050/76000 = $                5.86
Business = 1214100/760000 = $                1.60